# A Gross-Pitaevskii theory for an excitonic incompressible Bose solid

### v1

@@ Appendix A Derivation of Gross-Pitaevskii equation for an exciton Bose solid
We outline our generalized Gross-Pitaevskii formalism. Consider the wave function for a Bose solid [2] in the normal state,
|$\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=\displaystyle\prod_{m=1}^{N}\hat{c}_{m}^{\dagger}\ket{\Psi_{\mathrm{vac}}(\mathbf{r})}\ .$ (12)
|$\ket{\Psi_{\mathrm{vac}}(\mathbf{r})}$ is the vacuum, $\hat{c}_{m}^{\dagger}=\int d^{2}\mathbf{r}\,\phi_{m}^{\star}(\mathbf{r})\widehat{\psi}^{\dagger}(\mathbf{r})$, $\hat{c}_{m}=\int d^{2}\mathbf{r}\,\phi_{m}(\mathbf{r})\widehat{\psi}(\mathbf{r})$. The $\phi_{m}(\mathbf{r})$ are a set of identical functions, each centered on the different sites $\{m\}$ of a periodic lattice. They are degenerate and mutually orthogonal.
The effective Hamiltonian (Eq. (333) in the main manuscript) can be written,
^$\widehat{{\cal{H}}}^{\rm{eff}}=\sum_{i,j}\hat{c}_{i}^{\dagger}T_{ij}\hat{c}_{j}+\frac{1}{2}\!\!\sum_{i,i^{\prime},j,j^{\prime}}\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}^{\dagger}W_{i,i^{\prime},j,j^{\prime}}\hat{c}_{j}\hat{c}_{j^{\prime}}\ ,$ (13)
@@
The expectation energy functional for the state |$\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}$ is,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=$ $\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{m=1}^{N}\hat{c}_{m}\right)\sum_{i,j}\hat{c}_{i}^{\dagger}T_{ij}\hat{c}_{j}\left(\displaystyle\prod_{m^{\prime}=1}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)$ (16)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{m=1}^{N}\hat{c}_{m}\right)\frac{1}{2}\!\!\sum_{\begin{subarray}{c}i,i^{\prime},j,j^{\prime}\end{subarray}}\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}^{\dagger}W_{i,i^{\prime},j,j^{\prime}}\hat{c}_{j}\hat{c}_{j^{\prime}}\left(\displaystyle\prod_{m^{\prime}=1}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)\ .$
Note that in the interaction term of Eq. (16) the contributions in the summation $j\!=\!j^{\prime}$ and $i\!=\!i^{\prime}$ will give zero. This is because for the state |$\ket{\Psi_{\mathrm{BS}}}$ there is only one particle on each site $j$ , so $\left(\hat{c}_{j}\right)^{2}\ket{\Psi_{\mathrm{BS}}}=0$ and $\bra{\Psi_{\mathrm{BS}}}\left(\hat{c}_{j}^{\dagger}\right)^{2}\!=0$. We will see that this property is crucial for obtaining the boson solid. Then,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}$
$=\sum_{i,j}\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{\begin{subarray}{c}m=1\\ m\neq i\end{subarray}}^{N}\hat{c}_{m}\right)T_{ij}\left(\displaystyle\prod_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq j\end{subarray}}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)+\!\frac{1}{2}\!\!\sum_{\begin{subarray}{c}i,i^{\prime},j,j^{\prime}\\ i\neq i^{\prime},j\neq j^{\prime}\end{subarray}}\left(\bra{{\psi}_{\mathrm{vac}}(\mathbf{r})}\displaystyle\prod_{\begin{subarray}{c}m=1\\ m\neq i,i^{\prime}\end{subarray}}^{N}\hat{c}_{m}\right)W_{i,i^{\prime},j,j^{\prime}}\left(\displaystyle\prod_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq j,j^{\prime}\end{subarray}}^{N}\hat{c}_{m^{\prime}}^{\dagger}\ket{{\psi}_{\mathrm{vac}}(\mathbf{r})}\right)$
@@
Using completeness $\sum_{j}\phi_{j}^{\star}(\mathbf{r})\phi_{j}(\mathbf{r}^{\prime})\!=\!\delta(\mathbf{r}-\mathbf{r}^{\prime})$, we recognize that the first term in the interaction in Eq. (17) is zero since $t_{\text{XX}}(|\mathbf{r^{\prime}}-\mathbf{r}|\!=\!0)=0$, leaving,
$\bra{\Psi_{\mathrm{BS}}(\mathbf{r})}\widehat{{\cal{H}}}^{\rm{eff}}\ket{\Psi_{\mathrm{BS}}(\mathbf{r})}=\sum_{i}\int\,d^{2}\mathbf{r}\,\phi_{i}^{\star}(\mathbf{r})\left(-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\right)\phi_{i}(\mathbf{r})+\frac{1}{2}\sum_{\begin{subarray}{c}j,j^{\prime}\\ j\neq j^{\prime}\end{subarray}}\iint\!\!d^{2}\mathbf{r}\,d^{2}\mathbf{r}^{\prime}|\phi_{j^{\prime}}(\mathbf{r})|^{2}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{j}(\mathbf{r}^{\prime})|^{2}\ .$ (18)
@@
For an exciton quantum condensate, the macroscopic occupation of the condensate implies that a single-particle function, the order parameter, describes the many-body ground state. Instead of the many-body wave function for the normal state system (Eq. 12 that consists of a product of different single-particle wave functions each to describe an exciton on its site) the order parameter that we choose for the incompressible supersolid is of the form,
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\phi_{i}(\mathbf{r})\ ,$ (22)
for $N$ sites with $N$ particles. A system described by the order parameter ((22)) has off-diagonal long-range order in the presence of a non-zero overlap between adjacent sites, establishing phase coherence. The $\Psi(\mathbf{r})$ will have a maximum at every lattice site $i$, so additionally it also has diagonal long-range order due to the lattice symmetry. The combination of both off-diagonal and diagonal long-range order makes Eq. (22) a candidate for a supersolid order parameter.
For a supersolid there must exist phase locking between adjacent sites arising from mutual overlap of the $\phi_{i}(\mathbf{r})$ [19]. To monitor the overlap, we parametrize the order parameter, writing $\phi_{i}$ in the form,
$\phi_{i}(\mathbf{r})$ $=$ $\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})$ (23)
$\Psi(\mathbf{r})$ $=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left(\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{i})\right)\ .$ (24)
The background component $\sqrt{\rho_{B}}$ is uniform throughout the sample, and all the $G(\mathbf{r}-\mathbf{r}_{i})$ are the same function but localized on the different sites $i$ centered at $\mathbf{r}_{i}$.
Summing Eq. (21) over the sites $i$ and using Eq. (22) we obtain,
@@
$-\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\sum_{i}\phi_{i}(\mathbf{r})+\sum_{i}\!\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\sum_{i}\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\mu\!\sum_{i}\phi_{i}(\mathbf{r})$ (25)
$-\sqrt{N}\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\Psi(\mathbf{r})+\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\sum_{i}\!\!\int\!\!d^{2}\mathbf{r}^{\prime}\phi_{i}(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\sqrt{N}\mu\Psi(\mathbf{r})$ (26)
From Eq. (23), within the second term of Eq. (26) we can write,
$\sum_{j}|\phi_{j}(\mathbf{r})|^{2}$ $=$ $\sum_{j}\left|\frac{\sqrt{\rho_{B}}}{N}+G(\mathbf{r}-\mathbf{r}_{j})\right|^{2}=\left|\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}\sum_{j}G(\mathbf{r}-\mathbf{r}_{j})+\sum_{j}G^{2}(\mathbf{r}-\mathbf{r}_{j})\right|\ ,$ (27)
@@
When $G(\mathbf{r}-\mathbf{r}_{i})$ is localized within a single unit cell $i$, then for a value of $\mathbf{r}^{\prime}=\mathbf{r}_{0}$ located in the unit cell $i_{0}$, the summations over $j$ collapse to,
$\sum_{i}|\phi_{i}(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}G(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})+G^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ .$ (28)
Using Eq. (24),
$|\Psi(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})+\frac{1}{N}\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}-\mathbf{r}_{i^{\prime}})\ .$ (29)
The final term $\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}})$ in Eq. (29) is periodic since it is invariant with shifts in $\mathbf{r}\rightarrow\mathbf{r}+\mathbf{n}\cdot\mathbf{a}$, for $\mathbf{n}=\{n_{x},n_{y}\}$ any pair of integers. So in the summation over $i$, we may choose for each $i$ a different value for $\mathbf{n}_{i}$ such that $(\mathbf{r}_{i}+\mathbf{n}_{i}\cdot\mathbf{a})$ is always the unit cell containing $\mathbf{r}_{0}$. Using this construction, each term in the summation over $i$ becomes equal. Thus we may truncate the summation over $i$ to the single value $i=i_{0}$, multiplying by a factor $N$ to account for the $N$ equal terms,
$\sum_{i}G(\mathbf{r}_{0}-\mathbf{r}_{i})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}})$ $=$ $\sum_{i}\left\{G(\mathbf{r}_{0}-\mathbf{r}_{i}-\mathbf{n}_{i}\cdot\mathbf{a})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}}-\mathbf{n}_{i}\cdot\mathbf{a})\right\}$ (30)
$=$ $NG(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\sum_{i^{\prime}}G(\mathbf{r}_{0}-\mathbf{r}_{i^{\prime}}-\mathbf{n}_{i_{0}}\cdot\mathbf{a})$
$=$ $NG^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ ,$ (31)
since in the sum over $i^{\prime}$ in Eq. (30), only the term $(\mathbf{r}_{i^{\prime}}+\mathbf{n}_{i_{0}}\cdot\mathbf{a})=\mathbf{r}_{i_{0}}$ is non-zero. Equation (29) can thus be written,
$|\Psi(\mathbf{r}_{0})|^{2}$ $=$ $\frac{\rho_{B}}{N}+2\frac{\sqrt{\rho_{B}}}{N}G(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})+\frac{1}{N}NG^{2}(\mathbf{r}_{0}-\mathbf{r}_{i_{0}})\ ,$ (32)
and we conclude from Eqs. (28) and (32),
$|\Psi(\mathbf{r})|^{2}=\sum_{i}|\phi_{i}(\mathbf{r})|^{2}\ ,$ (33)
Hence in Eq. (26) the second term becomes,
$\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\psi_{j}(\mathbf{r}^{\prime})|^{2}=\sqrt{N}\!\int\!\!d^{2}\mathbf{r}^{\prime}\Psi(\mathbf{r})\ t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\Psi(\mathbf{r}^{\prime})|^{2}\ .$ (34)
In Eq. (26), the final term on the left-hand side has the same structure as the second term, but with the restriction $j=i$. We can accordingly use the same construction, but we must capture the index restriction $j=i$ by limiting the $\mathbf{r}^{\prime}$ integration to the unit cell containing $\mathbf{r}$. We use the notation ${\mathbf{r^{\prime}}\in{j_{\mathbf{r}}}}$. So we finally obtain,

### v2

@@ Appendix A Derivation of Gross-Pitaevskii equation for an exciton Bose solid
We outline here the development of our generalized Gross-Pitaevskii formalism.
Consider the wave function for a Bose solid [2] in the normal state,
[…]
The effective Hamiltonian (Eq. ($3$) in the main manuscript) can be written,
$\widehat{{\cal{H}}}^{\rm{eff}}=\sum_{i,j}\hat{c}_{i}^{\dagger}T_{ij}\hat{c}_{j}+\frac{1}{2}\!\!\sum_{i,i^{\prime},j,j^{\prime}}\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}^{\dagger}W_{i,i^{\prime},j,j^{\prime}}\hat{c}_{j}\hat{c}_{j^{\prime}}\ ,$ (11)
@@
[…]
@@
Using completeness $\sum_{j}\phi_{j}^{\star}(\mathbf{r})\phi_{j}(\mathbf{r}^{\prime})\!=\!\delta(\mathbf{r}-\mathbf{r}^{\prime})$, we recognize that, since $t_{\text{XX}}(|\mathbf{r^{\prime}}-\mathbf{r}|\!=\!0)=0$, the first interaction term in Eq. (15) is zero. This leaves,
[…]
@@
We turn now to the possibility of solidification coexisting with quantum coherence in a supersolid. Equation (10) for the many-body wave function of the exciton normal state solid consists of a product of different single-particle wave functions, each to describe an exciton on its site. In contrast, the many-body wave function for the exciton supersolid consists of a single function, the order parameter. The order parameter must possess off-diagonal long-range order, depending on a finite overlap between adjacent sites to establish phase coherence, and it must also possess diagonal long-range order due to the lattice symmetry. We propose the following form for the order parameter,
$\Psi(\mathbf{r})=\sqrt{\rho_{B}}+\sum_{i=1}^{N}\phi_{i}(\mathbf{r}-\mathbf{r}_{i})\ .$ (20)
for $N$ sites with $N$ particles. Each $\phi_{i}(\mathbf{r}-\mathbf{r}_{i})$ is the same function but localized on different sites $i$ and centered at $\mathbf{r}_{i}$. Since $\Psi(\mathbf{r})$ has a maximum at every lattice site $i$ it possesses the diagonal long-range order of the lattice symmetry. The other component $\sqrt{\rho_{B}}$ is uniform throughout the sample and ensures the phase locking between adjacent sites made possible from the mutual overlap of the sites. This is essential for the existence of a supersolid [19].
The combination of off-diagonal and diagonal long-range order makes Eq. (20) suitable as the order parameter for a supersolid.
Summing Eq. (19) over the sites $i$,
@@
$\frac{\hbar^{2}\nabla^{2}}{2M_{\text{X}}}\sum_{i}\phi_{i}(\mathbf{r})+\sum_{i}\phi_{i}(\mathbf{r})\!\int\!\!d^{2}\mathbf{r}^{\prime}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)\sum_{\begin{subarray}{c}j\end{subarray}}|\phi_{j}(\mathbf{r}^{\prime})|^{2}-\sum_{i}\phi_{i}(\mathbf{r})\!\int\!\!d^{2}\mathbf{r}^{\prime}t_{XX}(|\mathbf{r}^{\prime}\!\!-\!\mathbf{r}|)|\phi_{i}(\mathbf{r}^{\prime})|^{2}=\mu\!\sum_{i}\phi_{i}(\mathbf{r})$ (21)
@@
When the function $\phi_{i}(\mathbf{r}-\mathbf{r}_{i})$ is localized within a single unit cell $i$,